OpenStudy (anonymous):

The function f(x) changes value when x changes from x0 to x0 + dx. Find the approximation error . Round your answer, if appropriate. f(x) = x^2, x0 = 7, dx = 0.06 I forgot the steps of how to get started...can someone show me the first step please??

OpenStudy (anonymous):

Multiply out: \[(x_0+dx)^2\]

OpenStudy (anonymous):

The error is the last piece, the \((dx)^2\) piece.

OpenStudy (anonymous):

that's not one of the answer choices... i got .0036? is that right

OpenStudy (anonymous):

Did you substitute in the numbers?

OpenStudy (anonymous):

(7+.06)^2=49.84...

OpenStudy (anonymous):

Yes, \( (dx)^2=(0.06)^2\)

OpenStudy (anonymous):

ohhh okay. so (0.06)^2 is the answer?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

oh. thank you!